# The physical and chemical processes in protoplanetary disks: constraints on the composition of comets

### v3

@@ 3.3 Observations of ices in disks
Ices in disks have been observed when the disk is viewed close to edge-on, where ice in a flared outer disk absorbs thermal emission from hot dust in the inner disk, light scattered off the disk surface, or some combination thereof (Pontoppidan et al., 2005; Terada et al., 2007). Interpretation of such direct absorption spectroscopy is complicated by the complex geometry of the radiative transfer, making it difficult to establish unambiguously the location of the ices within the disk. Further, the signal-to-noise ratios of edge-on disk spectra have often been low due to the inherent faintness of edge-on disks at most infrared wavelengths. In many cases, part of the absorption may also be due to cold material in a remnant protostellar envelope, rather than in the disk itself. In the most unambiguous cases, it appears that the most abundant ices have relative concentrations consistent with those observed in dense clouds and protostellar envelopes (Aikawa et al., 2012), although more sensitive observations of rarer species, such as CO and CHOH, will likely be needed to establish evolutionary patterns (Ballering et al., 2021).
@@
Observations of the molecular inventory in the inner disk may be used to constrain the efficiency of radial transport of ices. Najita et al. (2013) found a trend between the mid-infrared HCN/HO flux ratio and dust disk mass as measured by ALMA. The proposed interpretation of this apparent link between inner and outer disk properties is that more massive disks may have experienced more efficient build-up of large, water-rich planetesimals, leaving less water ice available in dust grains that migrate inwards. In this case, the observed HCN/HO ratio is also a measure of the inner disk elemental C/O ratio. More recently, this scenario was supported by the observation of Banzatti et al. (2020), who found a similar relation of enhanced water emission for disks with small dust emission radii, suggesting that compact disks have experienced strong inward radial drift of water-ice rich dust, thereby enriching the inner disk with water vapor and decreasing the local C/O ratio (see also §4.2).
@@ 3.4.1 Isotope Fractionation
Since these reactions are exothermic, the backward reactions are inefficient at low temperatures ( several 10 K). The high D/H ratio (e.g. HD/H) propagates to other molecules via chemical reactions; e.g. HDO is formed by grain surface reactions of O atom or OH with D atom, whose abundance is enhanced by the dissociative recombination of HD. The high D/H ratio of the ocean suggests that water is supplied to Earth from cold regions, at least partially, and has been motivating the measurement of D/H ratios in comets. The D/H ratio of cometary water, , overlaps with that of ocean (Mumma and Charnley, 2011; Ceccarelli et al., 2014; Altwegg et al., 2015, Chapter by Biver et al.). Recent observations of higher DO/HDO ratio than HDO/HO both in comet 67P/C-G and protostellar cores indicate the inheritance of comet water from molecular clouds (Furuya et al., 2016, 2017; Altwegg et al., 2017), while the variation of HDO/HO ratio among comets indicates some reprocessing in disks.
@@
In protoplanetary disks, five deuterated molecules, DCO, ND, DCN, CD, and HDO are detected so far (e.g. Huang et al., 2017; Salinas et al., 2017; Cataldi et al., 2021; Öberg et al., 2021a; Loomis et al., 2020; Tobin et al., 2023) (Table 1). Disk averaged D/H ratio of the former three species are summarized in Table 2. It should be noted that the recombination timescale of ions is much shorter than the typical ages of disks (i.e. a few yr). The high D/H ratio of molecular ions is thus clear evidence of active deuterium fractionation in disks. The D/H ratio of NH is higher than that of HCO. It is consistent with the theoretical expectation that NH traces the cold ( K) layer near the midplane where CO is frozen out, which enhances the deuterium fractionation (Willacy, 2007; Cleeves et al., 2014; Aikawa et al., 2015, 2018).
@@
Recently, Tobin et al. (2023) detected HDO and HO in the disk of FU Ori star V883 ori, and derived the HDO/HO ratio to be . Since this ratio is comparable to those observed in the warm central protostellar cores, they conclude that water in the disk inherit from molecular clouds.
@@
Since temperature increases radially inwards in disks, the molecular D/H ratio is expected to decrease inwards, which is confirmed by the spatially resolved observations of DCN, DCO and ND (Cataldi et al., 2021; Öberg et al., 2021a). The radial distribution of the molecular D/H ratio could also probe the deuteration pathways of each molecular species. The decline of D/H ratio towards the warm central region is expected to be less steep if the molecule is formed mainly via hydrocarbons rather than H, since the endothermicity of reaction (6) is higher than that of (5) (Öberg et al., 2012). Cataldi et al. (2021) derived the DCN/HCN ratio as a function of the excitation temperature of HCN (i.e. proxy of gas temperature), which indicates the significant contribution of reaction (6). They also found that the DCN/HCN ratio becomes comparable to the value measured in Hale-Bopp, (Meier et al., 1998; Crovisier et al., 2004), in the regions of K.
@@
The N/N ratio of HCN, CN, NH has been measured in comets. The ratio is lower than the elemental abundance in the local interstellar medium () (e.g. Ritchey et al., 2015) and the protosolar value (441) (Marty et al., 2011). For example, the average CN/CN ratio over 20 comets is (Manfroid et al., 2009). The N enrichment could be due to exchange reactions and/or selective photodissociation of N. The former would be less effective than previously thought, since some key exchange reactions are found to have activation barriers (Roueff et al., 2015). In the latter mechanism, NN is photodissociated in deeper layers of a disk or molecular cloud than the major isotope, producing excess N atoms, which are incorporated to other molecules such as CN (Liang et al., 2007; Heays et al., 2014; Furuya et al., 2017). Visser et al. (2018) calculated N/N fractionation in disk models to find that the fractionation is fully dominated by selective photodissociation of N (see also Lee et al., 2021).
@@
Guzmán et al. (2017) observed HCN and HCN to find HCN/HCN ratio of 80-160 in five disks, which are roughly consistent with the values in comets. In the disk of TW Hya (Table 3), Hily-Blant et al. (2019) found that the HCN/HCN ratio increases with radius, which is consistent with the fractionation due to selective photodissociation (see also Guzmán et al., 2017). The ratio of CN/CN, on the other hand, is (Hily-Blant et al., 2019). The different N/N ratio in HCN and CN could be a natural outcome of the vertical gradient of molecular abundance; i.e. CN has its abundance peak in the upper layer than that of HCN (Cazzoletti et al., 2018b; Teague and Loomis, 2020) (§3.1).
@@
Carbon and oxygen isotope ratios in comets are mostly consistent with the Solar abundance, with the notable exception of some molecules in 67P/C-G (see the chapter by Biver et al. for more details). On the theoretical side, selective photodissociation of CO, combined with isotope exchange reaction of C + CO CO + C, could induce fractionation of C/C and O/O/O (Yurimoto and Kuramoto, 2004; Lyons and Young, 2005). Miotello et al. (2014) calculated the isotope fractionation of CO in disk models to show that the abundance and thus the line flux of CO and CO could be significantly reduced by the selective photodissociation, especially in disks with grain growth. Observational confirmation is challenging, since the lines of the major isotope of CO are optically thick, while the rare isotope lines are very weak. Smith et al. (2009) found that the CO/CO and CO/CO ratios are lower than the elemental abundances in the disk of VV CrA by observing absorption lines in the near-infrared with the binary star as a possible light source. As for carbon, Hily-Blant et al. (2019) found that the HCN/HCN ratio is mostly flat with at au in the disk of TW Hya. Yoshida et al. (2022), on the other hand, derived CO/CO at au, which indicates the fractionation via the isotope fractionation (C + CO) in the gas of high C/O () elemental abundance. Recently, Furuya et al. (2022b) detected HCO emission in TW Hya. Combining this detection with previous HCO observation and disk chemical model, CO/CO ratio is estimated to be consistent with the elemental abundance ratio in the local ISM.
@@ 3.4.2 Spin Temperature
Hydrogen atoms have nuclear spin angular momentum, and molecules that contain two or more H atoms in symmetrical locations have quantum states distinguished by the total spin angular momentum (). In the case of HO, for example, for ortho (-HO) and for para (-HO). Spontaneous conversion of spin state (e.g. ortho to para) is strongly forbidden by the quantum mechanical selection rules. The rotational energy levels for each spin state form a discrete ladder, and the energy of the lowest-lying level is lower for -HO than -HO by 34.2 K. The relative population of the ortho-to-para ratio (OPR) of HO is described as
@@
The OPR of HO, NH, and some other hydrated molecules have been measured in comets, assuming that is equal to the temperature of the formation site of the molecule (e.g. Mumma et al., 1987; Mumma and Charnley, 2011; Shinnaka et al., 2020). The derived spin temperature range is K with the typical value of K. For HO, the corresponding OPR is (Mumma and Charnley, 2011).
@@
The interpretation of as the temperature of the molecular formation site, however, turned out to be too simplistic. A laboratory experiment by Hama et al. (2018) showed that the OPR of HO desorbed from ice is 3 regardless of the temperature of the ice formation (see also Hama et al., 2016). While the energy difference of the lowest-lying state between ortho and para is 34.2 K in the gas phase, the rotation motion is constrained and the energy difference becomes significantly small in the ice phase. Then the magnetic dipole interaction of a proton with neighboring protons enables rapid spin conversion (Limbach et al., 2006; Hama et al., 2018).
The lower OPR than the statistical value observed in comets and outer protoplanetary disks would thus be due to the gas-phase reactions; the spin state changes via proton exchange reactions such as -HO + H -HO + H. The cold water vapor in protoplanetary disks is much less abundant compared with the value expected from the desorption of ice (§3.1). The gas-phase formation of HO could thus play a role, while it is much less efficient than the grain-surface reactions at K. As for comets, the effect of sublimation processes and gas-phase reactions in the coma needs to be considered for the interpretation of the observed OPR ratios (Hama and Watanabe, 2013; Hama et al., 2016)
@@ 4.1 From molecular clouds to disks
It has long been discussed whether the molecular composition of volatiles in comets inherits ices in molecular clouds or is reset in the Solar nebula (e.g. Mumma and Charnley, 2011). To pursue this question, we need to investigate not only Class II disks but also the disk formation stage (i.e. Class 0-I). Studies of forming disks have seen significant progress in the last decade.
@@
On the theoretical side, disk formation has been studied by hydrodynamics simulations starting from the collapse of the parental cloud core, considering the effect of radiation transfer and magnetic fields to derive the physical structures (e.g. size, mass, temperatures) of the disks as a function of time (e.g. Machida et al., 2011; Tsukamoto et al., 2015; Hennebelle et al., 2016; Tsukamoto et al., 2023). The molecular evolution of gas and ice is investigated by calculating the chemical reaction network along the gas flow from the cloud core to the disk, considering the temporal variation of density, temperature, and UV radiation flux along the trajectories (e.g. Visser et al., 2009, 2011; Yoneda et al., 2016; Drozdovskaya et al., 2016; Furuya et al., 2017; Aikawa et al., 2020). Stable abundant ices such as HO and CO are delivered to the disk without significant alteration, unless the trajectory is close to the outflow cavity where photodissociation is effective. [Outflows create a cavity (a pair of cone-shaped low-density regions) in the envelope. The wall of the outflow cavity could be irradiated by the UV radiation from the central object.] The abundance of COMs, on the other hand, could increase in the ice mantle as the thermal diffusion and reactions are activated in warm circumstellar conditions (e.g. Garrod and Herbst, 2006; Garrod et al., 2022).
@@
Observations of forming disks are more complicated than those of Class II, since the emission from the disk needs to be distinguished from the infalling envelope and outflow. Careful analysis of spatial distributions and velocity structures shows that emission lines of rare isotopes of CO (with optical depth lower than their normal isotope), HCO, and CS tend to probe the disk, while SO traces week accretion shock onto the disk (Sakai et al., 2014; Aso et al., 2017; Oya et al., 2018; Harsono et al., 2021; Tychoniec et al., 2021; Garufi et al., 2021). The vertical stratification of molecules similar to that in Class II disks (§3.1) is also found (Podio et al., 2020). van’t Hoff et al. (2020) show that the young disks tend to be warmer than Class II disks, with the CO snowline at several tens of au, possibly indicating a higher accretion rate (§2.2). Disks with warped structures are also found (e.g. Sakai et al., 2019; Sai et al., 2020). Recent observations found large-scale (thousands of au) non-axisymmetric steamers onto young disks (Yen et al., 2014; Pineda et al., 2020; Garufi et al., 2022). Such a flow provides chemically fresh interstellar material (e.g. carbon chains) and angular momentum that could perturb the disk structure to enhance mixing.

### v4

@@ 3.3 Observations of ices in disks
Ices in disks have been observed when the disk is viewed close to edge-on, where ice in a flared outer disk absorbs thermal emission from hot dust in the inner disk, light scattered off the disk surface, or some combination thereof (Pontoppidan et al., 2005; Terada et al., 2007). Interpretation of such direct absorption spectroscopy is complicated by the complex geometry of the radiative transfer, making it difficult to establish unambiguously the location of the ices within the disk. Further, the signal-to-noise ratios of edge-on disk spectra have often been low due to the inherent faintness of edge-on disks at most infrared wavelengths. In many cases, part of the absorption may also be due to cold material in a remnant protostellar envelope, rather than in the disk itself. In the most unambiguous cases, it appears that the most abundant ices have relative concentrations consistent with those observed in dense clouds and protostellar envelopes (Aikawa et al., 2012), although more sensitive observations of rarer species, such as CO2 and CH3OH, will likely be needed to establish evolutionary patterns (Ballering et al., 2021).
@@
Observations of the molecular inventory in the inner disk may be used to constrain the efficiency of radial transport of ices. Najita et al. (2013) found a trend between the mid-infrared HCN/H2O flux ratio and dust disk mass as measured by ALMA. The proposed interpretation of this apparent link between inner and outer disk properties is that more massive disks may have experienced more efficient build-up of large, water-rich planetesimals, leaving less water ice available in dust grains that migrate inwards. In this case, the observed HCN/H2O ratio is also a measure of the inner disk elemental C/O ratio. More recently, this scenario was supported by the observation of Banzatti et al. (2020), who found a similar relation of enhanced water emission for disks with small dust emission radii, suggesting that compact disks have experienced strong inward radial drift of water-ice rich dust, thereby enriching the inner disk with water vapor and decreasing the local C/O ratio (see also §4.2).
@@ 3.4.1 Isotope Fractionation
Since these reactions are exothermic, the backward reactions are inefficient at low temperatures ( several 10 K). The high D/H ratio (e.g. H2D+/H) propagates to other molecules via chemical reactions; e.g. HDO is formed by grain surface reactions of O atom or OH with D atom, whose abundance is enhanced by the dissociative recombination of H2D+. The high D/H ratio of the ocean suggests that water is supplied to Earth from cold regions, at least partially, and has been motivating the measurement of D/H ratios in comets. The D/H ratio of cometary water, , overlaps with that of ocean (Mumma and Charnley, 2011; Ceccarelli et al., 2014; Altwegg et al., 2015, Chapter by Biver et al.). Recent observations of higher D2O/HDO ratio than HDO/H2O both in comet 67P/C-G and protostellar cores indicate the inheritance of comet water from molecular clouds (Furuya et al., 2016, 2017; Altwegg et al., 2017), while the variation of HDO/H2O ratio among comets indicates some reprocessing in disks.
@@
In protoplanetary disks, five deuterated molecules, DCO+, N2D+, DCN, C2D, and HDO are detected so far (e.g. Huang et al., 2017; Salinas et al., 2017; Cataldi et al., 2021; Öberg et al., 2021a; Loomis et al., 2020; Tobin et al., 2023) (Table 1). Disk averaged D/H ratio of the former three species are summarized in Table 2. It should be noted that the recombination timescale of ions is much shorter than the typical ages of disks (i.e. a few yr). The high D/H ratio of molecular ions is thus clear evidence of active deuterium fractionation in disks. The D/H ratio of N2H+ is higher than that of HCO+. It is consistent with the theoretical expectation that N2H+ traces the cold ( K) layer near the midplane where CO is frozen out, which enhances the deuterium fractionation (Willacy, 2007; Cleeves et al., 2014; Aikawa et al., 2015, 2018).
@@
Recently, Tobin et al. (2023) detected HDO and HO in the disk of FU Ori star V883 ori, and derived the HDO/H2O ratio to be . Since this ratio is comparable to those observed in the warm central protostellar cores, they conclude that water in the disk inherit from molecular clouds.
@@
Since temperature increases radially inwards in disks, the molecular D/H ratio is expected to decrease inwards, which is confirmed by the spatially resolved observations of DCN, DCO+ and N2D+ (Cataldi et al., 2021; Öberg et al., 2021a). The observed peak D/H ratios and their radial positions are summarized in Table 2 (see also Figure 5). The radial distribution of the molecular D/H ratio could also probe the deuteration pathways of each molecular species. The decline of D/H ratio towards the warm central region is expected to be less steep if the molecule is formed mainly via hydrocarbons rather than H, since the endothermicity of reaction (6) is higher than that of (5) (Öberg et al., 2012). Cataldi et al. (2021) derived the DCN/HCN ratio as a function of the excitation temperature of HCN (i.e. proxy of gas temperature), which indicates the significant contribution of reaction (6). They also found that the DCN/HCN ratio becomes comparable to the value measured in Hale-Bopp, (Meier et al., 1998; Crovisier et al., 2004), in the regions of K.
@@
The 14N/15N ratio of HCN, CN, NH2 has been measured in comets. The ratio is lower than the elemental abundance in the local interstellar medium () (e.g. Ritchey et al., 2015) and the protosolar value (441) (Marty et al., 2011). For example, the average C14N/C15N ratio over 20 comets is (Manfroid et al., 2009). The 15N enrichment could be due to exchange reactions and/or selective photodissociation of N2. The former would be less effective than previously thought, since some key exchange reactions are found to have activation barriers (Roueff et al., 2015). In the latter mechanism, 14N15N is photodissociated in deeper layers of a disk or molecular cloud than the major isotope, producing excess 15N atoms, which are incorporated to other molecules such as CN (Liang et al., 2007; Heays et al., 2014; Furuya et al., 2017). Visser et al. (2018) calculated 14N/15N fractionation in disk models to find that the fractionation is fully dominated by selective photodissociation of N2 (see also Lee et al., 2021).
@@
Guzmán et al. (2017) observed H13CN and HC15N to find HCN/HC15N ratio of 80-160 in five disks, which are roughly consistent with the values in comets. In the disk of TW Hya (Table 3), Hily-Blant et al. (2019) found that the HCN/HC15N ratio increases with radius, which is consistent with the fractionation due to selective photodissociation (see also Guzmán et al., 2017). The ratio of CN/C15N, on the other hand, is (Hily-Blant et al., 2019). The different 14N/15N ratio in HCN and CN could be a natural outcome of the vertical gradient of molecular abundance; i.e. CN has its abundance peak in the upper layer than that of HCN (Cazzoletti et al., 2018b; Teague and Loomis, 2020) (§3.1).
@@
Carbon and oxygen isotope ratios in comets are mostly consistent with the Solar abundance, with the notable exception of some molecules in 67P/C-G (see the chapter by Biver et al. for more details). On the theoretical side, selective photodissociation of CO, combined with isotope exchange reaction of 13C+ + CO 13CO + C+, could induce fractionation of 12C/13C and 16O/17O/18O (Yurimoto and Kuramoto, 2004; Lyons and Young, 2005). Miotello et al. (2014) calculated the isotope fractionation of CO in disk models to show that the abundance and thus the line flux of C18O and C17O could be significantly reduced by the selective photodissociation, especially in disks with grain growth. Observational confirmation is challenging, since the lines of the major isotope of CO are optically thick, while the rare isotope lines are very weak. Smith et al. (2009) found that the C18O/CO and C17O/CO ratios are lower than the elemental abundances in the disk of VV CrA by observing absorption lines in the near-infrared with the binary star as a possible light source. As for carbon, Hily-Blant et al. (2019) found that the HCN/H13CN ratio is mostly flat with at au in the disk of TW Hya. Yoshida et al. (2022), on the other hand, derived 12CO/13CO at au, which indicates the fractionation via the isotope fractionation (C+ + CO) in the gas of high C/O () elemental abundance. Furuya et al. (2022b) detected HC18O+ emission in TW Hya. Combining this detection with previous H13CO+ observation and disk chemical model, 13CO/C18O ratio is estimated to be consistent with the elemental abundance ratio in the local ISM. More recently, Yamato et al. (2024b) found that the 12C/13C ratios of freshly sublimated complex organic molecules in V883 Ori is low () (section 4.2).
@@ 3.4.2 Spin Temperature
Hydrogen atoms have nuclear spin angular momentum, and molecules that contain two or more H atoms in symmetrical locations have quantum states distinguished by the total spin angular momentum (). In the case of H2O, for example, for ortho (-H2O) and for para (-H2O). Spontaneous conversion of spin state (e.g. ortho to para) is strongly forbidden by the quantum mechanical selection rules. The rotational energy levels for each spin state form a discrete ladder, and the energy of the lowest-lying level is lower for -H2O than -H2O by 34.2 K. The relative population of the ortho-to-para ratio (OPR) of H2O is described as
@@
The OPR of H2O, NH3, and some other hydrated molecules have been measured in comets, assuming that is equal to the temperature of the formation site of the molecule (e.g. Mumma et al., 1987; Mumma and Charnley, 2011; Shinnaka et al., 2020). The derived spin temperature range is K with the typical value of K. For H2O, the corresponding OPR is (Mumma and Charnley, 2011).
@@
The interpretation of as the temperature of the molecular formation site, however, turned out to be too simplistic. A laboratory experiment by Hama et al. (2018) showed that the OPR of H2O desorbed from ice is 3 regardless of the temperature of the ice formation (see also Hama et al., 2016). While the energy difference of the lowest-lying state between ortho and para is 34.2 K in the gas phase, the rotation motion is constrained and the energy difference becomes significantly small in the ice phase. Then the magnetic dipole interaction of a proton with neighboring protons enables rapid spin conversion (Limbach et al., 2006; Hama et al., 2018).
The lower OPR than the statistical value observed in comets and outer protoplanetary disks would thus be due to the gas-phase reactions; the spin state changes via proton exchange reactions such as -H2O + H+ -H2O + H+. The cold water vapor in protoplanetary disks is much less abundant compared with the value expected from the desorption of ice (§3.1). The gas-phase formation of H2O could thus play a role, while it is much less efficient than the grain-surface reactions at K. As for comets, the effect of sublimation processes and gas-phase reactions in the coma needs to be considered for the interpretation of the observed OPR ratios (Hama and Watanabe, 2013; Hama et al., 2016)
@@ 4.1 From molecular clouds to disks
It has long been discussed whether the molecular composition of volatiles in comets inherits ices in molecular clouds or is reset in the Solar nebula (e.g. Mumma and Charnley, 2011). To pursue this question, we need to investigate not only Class II disks but also the disk formation stage (i.e. Class 0-I) (section 1). Studies of forming disks have seen significant progress in the last decade.
@@
On the theoretical side, disk formation has been studied by hydrodynamics simulations starting from the collapse of the parental cloud core, considering the effect of radiation transfer and magnetic fields to derive the physical structures (e.g. size, mass, temperatures) of the disks as a function of time (e.g. Machida et al., 2011; Tsukamoto et al., 2015; Hennebelle et al., 2016; Tsukamoto et al., 2023). The molecular evolution of gas and ice is investigated by calculating the chemical reaction network along the gas flow from the cloud core to the disk, considering the temporal variation of density, temperature, and UV radiation flux along the trajectories (e.g. Visser et al., 2009, 2011; Yoneda et al., 2016; Drozdovskaya et al., 2016; Furuya et al., 2017; Aikawa et al., 2020). Stable abundant ices such as H2O and CO2 are delivered to the disk without significant alteration, unless the trajectory is close to the outflow cavity where photodissociation is effective. [Outflows create a cavity (a pair of cone-shaped low-density regions) in the envelope. The wall of the outflow cavity could be irradiated by the UV radiation from the central object.] The abundance of COMs, on the other hand, could increase in the ice mantle as the thermal diffusion and reactions are activated in warm circumstellar conditions (e.g. Garrod and Herbst, 2006; Garrod et al., 2022).
@@
Observations of forming disks are more complicated than those of Class II, since the emission from the disk needs to be distinguished from the infalling envelope and outflow. Careful analysis of spatial distributions and velocity structures shows that emission lines of rare isotopes of CO (with optical depth lower than their normal isotope), H2CO, and CS tend to probe the disk, while SO traces week accretion shock onto the disk (Sakai et al., 2014; Aso et al., 2017; Oya et al., 2018; Harsono et al., 2021; Tychoniec et al., 2021; Garufi et al., 2021). The vertical stratification of molecules similar to that in Class II disks (§3.1) is also found (Podio et al., 2020). van’t Hoff et al. (2020) show that the young disks tend to be warmer than Class II disks, with the CO snowline at several tens of au, possibly indicating a higher accretion rate (§2.2). Disks with warped structures are also found (e.g. Sakai et al., 2019; Sai et al., 2020). Recent observations found large-scale (thousands of au) non-axisymmetric steamers onto young disks (Yen et al., 2014; Pineda et al., 2020; Garufi et al., 2022). Such a flow provides chemically fresh interstellar material (e.g. carbon chains) and angular momentum that could perturb the disk structure to enhance mixing.